### v1

@@ 5.4 Saliency Maps
To better understand the influence of the input features on our final predictions and also the difference our proposed method makes on how the networks determine what features to prioritize for prediction, we utilized a saliency map technique Simonyan et al. (2013) to visualize the importance of model weights on our input time axis and features. According to the predicted class $c$, for example, the decision-making process of the model can be represented as $S_{c}(I)=w_{c}^{T}I+b_{c}$, where $I,S_{c}(I),w,b$ are the model input, output, weights, and bias, respectively. The purpose of this method is to calculate the model weights on the input layer by gradients, for example, $w=\frac{\partial S_{c}}{\partial I}|_{I_{0}}$. The calculated $w$ represents the model saliency associated with the input layer. In this study, we fetched the saliency maps for all the test samples and calculated the average saliency map to develop the general intuition of important features and time steps.
We calculated the feature importance for our experiments according to the formulas in this section and plotted them in the figures in the results section.
@@ 6.1 ADULT Dataset
The initial analysis before running a model showed that all 3 protected labels biased the data. For example, for the age label, the DIR score was above 2 for both the training and testing data; for sex, it was approximately 0.36 for both, and for race, it was approximately 0.47 for both. After running the base model, this trend continued, where the model introduced even more bias by age, making the DIR above 3 and below 0.1 for sex, At the same time, for race, it stayed similar to the initial analysis with a score of approximately 0.5.
Since our analyses on the baseline model showed bias by all 3 chosen protected labels, we experimented with our proposed method on all 3 protected labels. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figures 4. To give a better, more linear visual representation of our results, we calculated the absolute value of [$1-$DIR Score] for the models to plot against the accuracy. The lines on the plots represent the Pareto fronts which contain the models that performed best to choose from. For all of the experiments on this dataset, we opted for the models that provided us the best fairness each time due to the insignificant differences in performance (accuracy) between the models (they all fall between 0.77 and 0.8). Therefore, we chose the models enclosed in the green circles to represent our final models.
@@
We also utilized the reweighing technique to improve fairness in these models as a comparison to our proposed method. We compared our final performance results of the base model, reweighing, and proposed method in Figure LABEL:adult_di_acc_plot. The reweighing technique slightly improved the fairness scores, with minimal impact on the performance scores of the models. For the protected age label, reweighing improved the DIR score by 0.84, for sex by 0.159, and for race by 0.0067 which we deemed not significant. The chosen model from our proposed method performed better than reweighing in terms of fairness, with an improvement in the DIR score of the age label by 1.28, sex by 0.32, and 0.042 for the race label. However, this improvement in fairness came at a slight performance cost, where the proposed method caused an average decrease of 1.7% in accuracy for all 3 labels. The reweighing method had an even lower performance cost, improving the accuracy score for the sex label. These results are visually represented in Figure 5 where the markers in blue represent the baseline models, red represents models implemented with reweighing technique, and green represents the models produced by our proposed method. In this figure, the x-axis represents an absolute value of the DIR score subtracted from 1 for an easier representation of our results.
@@
After calculating the feature importance for the baseline model and the models produced by our experiments for age, race, and sex labels, the scores we obtained showed us that our method redistributed weights to features less relevant to the protected label. This change is demonstrated in Figure 6. These figures show that the workclass feature holds the most weight across the board for all labels and remains equally important after the proposed method implementation. On the other hand, the relationship feature holds no weight in any of the prediction tasks across all experiments. The changes in importance are hence demonstrated more on the 3 remaining features: education, marital status, and occupation. Furthermore, the relative importance of these features changes across the experiments according to the demographic label we tested our proposed method on. For example, for the age label, the least important features are occupation, marital status, and education in that order, and the proposed method implemented with age shifted weight importance to these features. This trend is observed for the other labels of race and sex.
@@ 6.2 MIMIC-III Dataset
The baseline model for IHM prediction using the MIMIC-III dataset resulted in Accuracy = 0.866, AUROC = 0.603, and AUPRC = 0.181.
Our analysis of the MIMIC-III data showed that the baseline model is only biased by marital status out of the 4 protected labels we analyzed, with a DIR score of 1.3. Regarding age, gender, and insurance labels, the baseline model showed unsubstantial bias where the DIRs calculated for each fell within the acceptable range of 0.8-1.2, as shown in Table 3. The differences in false negative scores are also significantly low across all the protected labels.
@@
Since our analyses showed bias by marital status, we experimented with our proposed method using our target label of IHM and marital status. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figure 7 for DIR plotted against the chosen metric AUROC. To provide a better, more linear visual representation of our results, we calculated the absolute value of [$1-$DIR Score] for the models to plot against the accuracy as we did in 6.1. The lines on the plots represent the Pareto fronts which contain the models that performed best to choose from for each of the performance scores. Our proposed method improved the 3 performance scores for IHM prediction across the 4 protected labels. The general accuracy for the models on the Pareto front showed an average improvement of 1.74% from the baseline model. The AUROC score improved by an average of 0.20, and the AUPRC score by 0.197. In this experiment, we focused on prioritizing the AUROC score, as mentioned in section 5.1 because AUROC is prioritized for this dataset in previous studies Harutyunyan et al. (2019), and the accuracy variability for the models in the Pareto front is low (approximately 0.65%). We selected our final model from the AUROC vs. DI score Pareto plot (Figure 7). From this plot, the 2 circled models represent the optimal solutions from our method. The performance scores for these 2 models are represented in Table 4. We chose model 40 as our final solution since it provided an almost equal improvement in DIR as model 23, with a more significant improvement in AUROC, AUPRC, and accuracy.
Implementing the reweighing bias mitigation technique showed similar prediction performance when compared to our proposed method (as seen in Table 4), improving the AUROC, AUPRC, and accuracy slightly less than the proposed method when compared to the baseline. Reweighing improved the DIR score of the baseline model by 0.12 but did not perform as well as our proposed method. Regarding the difference in false negative scores, the reweighing technique performed worse than the baseline.
@@
Score Model 23 Model 40 Reweighing Baseline Model
Accuracy 0.879 0.883 0.883 0.866
AUROC 0.771 0.797 0.777 0.603
AUPRC 0.354 0.367 0.332 0.181
DI Ratio 0.99 1.043 1.188 1.308
Diff in FN 0.0199 0.005 -0.05 0.005
Figure 7: MIMIC-III: Pareto Front Plot for AUROC vs DI Ratio Score
@@ 6.3 SNAPSHOT Dataset
After performing the initial set of 32 fairness analyses on the SNAPSHOT data, our results showed that the data itself did not contain substantial bias by any of the 8 protected labels we tested by. The scatter plot in Figure 8 shows that all scores fall within the acceptable range of 1.2 and 0.8 for the DIR score, except for the experiment with Race as a protected label and evening-sad-happy as the target label (specified by the red circle on the figure). For the 4 target labels, morning-sad-happy, morning-stressed-calm, evening-sad-happy, and evening-stressed calm, the F1 performance scores fell from approximately 0.7 to 0.8.
@@
We ran our proposed method on Race as a protected label and evening-sad-happy as the target label and obtained accuracy scores in the range of 0.76 to 0.77 and DIR scores between 0.80 and 0.88 for the model ensemble produced. We chose the model with the best DIR score of 0.882 as shown in Figure 9. This is an improvement of approximately 0.1 in fairness from the baseline model with a DIR score of 0.78. In terms of performance, this method did not affect the F1 score when compared to the baseline model. Both have F1 scores of approximately 0.76. Using the reweighing technique, we obtained a perfect DIR score of 1, but that came at a high cost of performance, with the F1 score dropping to 0.39, a loss of 37% in performance.
@@ 6.3.1 Data Manipulation on SNAPSHOT
After analyzing the SNAPSHOT baseline model for bias and our results showed that the data was only biased by the Race label for evening-sad-happy label prediction; we performed some data manipulation to create a more significant imbalance for the gender label in order to determine if this would decrease the fairness performance and to test our proposed method on it. Originally, the data contained samples from 181 male and 132 female participants. There was a total of 4030 train and 1866 test samples, and we removed the samples from 75 participants that identified as female from the training data, making it a total of 2660 samples for training.We removed participants with 14 or more samples in the train data (experiment S-1). We also experimented with removing samples from participants with 5 or more samples in the train data (experiment S-2) and those with 30 or more samples (experiment S-3), leaving 2349 and 3685 samples, respectively. The motivation is to induce an artificial data bias, otherwise known as a negative legacy, into our model by creating a higher imbalance in the representation of male and female participant samples in our training data. With this manipulation, we expect our model to not generalize well on the test data that still contains a comparable number of female samples as hypothesized by Gorrostieta et al. (2019), thereby inducing some bias in its predictions.
After running the base model for all 4 target labels and analyzing fairness for all 3 experiment settings, we obtained similar performance scores. The fairness scores for experiments S-1 and S-2 were similar, with only the DIR score of the target label of evening-sad-happy occurring outside of the acceptable bounds of 0.8-1.2. In contrast, the fairness scores for experiment S-3 did not change significantly from the original SNAPSHOT experiment with no data manipulation. The F1 and fairness metric scores for experiment S-2 are listed in Table 5. As seen in the table, the DIR score of the target label of evening-sad-happy (the gray highlighted cell) fell outside the acceptable bounds of 0.8 and 1.2. All other fairness scores fell within these acceptable bounds.
@@
After running our proposed bias mitigation algorithm on the label combination of evening-sad-happy and gender on the manipulated data, we obtained an ensemble of models whose performance and DIR scores are plotted in Figure 10. On average, the models had an F1 score of 0.74 which is a 5.8% drop in performance from the baseline evening-sad-happy model. The average DIR score of the 9 models on the Pareto front is 1.034, which is well within the acceptable bounds of 0.8-1.2 and an overall improvement of 0.257 from the base DI ratio score of 1.291. We chose models 89 and 43 (green and red circles in Figure 10) as potential final models. Model 89 has the best DI ratio score of 1.001, which is almost the ideal score of 1, but it comes with a diminished F1 score of 0.722, a 7.6% drop in performance. It differs in False Positive and False Negative rates scores of -0.082 and -0.084, respectively. Model 43 has a DI ratio score of 0.995, with slightly better performance (0.733) than model 89.
@@ 7.1 Fairness and Performance
With the current re-occurring limitations that bias and fairness research in machine learning face, our study aims to develop an effective, easily transferable bias mitigation technique, with an option of determining the tradeoff regarding fairness and performance of the model. Our proposed method improved fairness when tested on models developed for three datasets, ADULT, MIMIC-III, and SNAPSHOT, with minimal depreciation in performance and improvement in the case of MIMIC-III. Compared to the standard bias mitigation technique of reweighing, our proposed method performed better in significantly improving fairness for the ADULT and MIMIC-III experiments and slightly improved fairness for the SNAPSHOT dataset, but at no performance cost, while reweighing diminishing the F1 score by 37%.
@@
From our experiments in predicting income with the ADULT dataset, our analyses showed us that the baseline model prediction of income was biased based on age, sex, and race, which corresponds with analyses from prior works (Le Quy et al. (2022); Lahoti et al. (2020); Kearns et al. (2019); Friedler et al. (2019); Zafar et al. (2017)). Implementing our proposed method provided a distribution of models that improved the DIR scores to within the ideal threshold of 0.8 and 1.2. This method performed better than the reweighing technique in terms of fairness, leading to an average of 1.7% decrease in accuracy for all three protected labels.
The LSTM model for IHM prediction using the MIMIC-III dataset was biased by marital status, not by age, gender, or insurance type. These results partially correlated with the fairness analysis performed in Röösli et al. (2022), where they found that gender was unaffected in terms of model fairness through their classification parity assessment Röösli et al. (2022). In their work, they did not consider marital status as a possible measure of fairness. In this work, we include marital status as a criterion due to various research showing evidence of the possible influence of marital status on one’s health. Dhindsa et al. (2020) showed that unmarried patients, including those who are divorced, separated, widowed, or never married have an increased rate of adverse cardiovascular events when compared to their married counterparts, and Johnson et al. (2000) found that non-married individuals show elevated rates of death compared to married persons, from the National Longitudinal Mortality Study (NLMS).
After finding that the model was biased by marital status, we implemented our bias-reducing technique, which improved the fairness to an almost-optimal level with a DIR score of 1.04 and a significantly low difference in False Negative scores between the privileged and unprivileged groups. In the process of doing this, our method also improved performance scores, raising the accuracy by 1.7%, AUROC score by 19.4%, and AUPRC by 18.6%. Compared to the reweighing technique, our method improved both fairness and prediction performances. We attribute this improvement in performance to a reason mentioned in prior research. As shown by the test for fairness-utility tradeoff conducted in Deho et al. (2022), ensuring fairness does not always lead to a drop in utility. Instead, it could lead to enhanced performance under specific circumstances.
@@
For SNAPSHOT, we analyzed baseline models that predicted 4 self-reported mood-related labels to determine if they were biased by 3 demographic labels, namely gender, race, ethnicity, and 5 personality-based labels, namely openness, conscientiousness, extraversion, agreeableness, and neuroticism. These analyses determined that the DIR scores from all our experiments were within the chosen acceptable range of 0.8 to 1.2, except for the model predicting evening-happy-sad when analyzed by race. Implementing our proposed method improved the DIR score to 0.88, an improvement of 0.1 from that of the baseline method. In terms of performance, our proposed method had little effect in this case. Implementing the reweighing technique provided a perfect DIR score of 1 but a considerable performance cost of 37%.
@@
As our initial analyses on the SNAPSHOT dataset baseline models did not show substantial bias by the 8 protected labels we used, we manipulated data by removing samples from 75 participants identified as female from the training data to determine the performance of our proposed method further. After running the baseline model on this altered data and analyzing, it showed bias by gender for the evening-sad-happy label with a DIR score of 1.29, slightly above our chosen acceptable range. Running our proposed technique resulted in an average, almost ideal DIR score of 1.034, an average 5.8% drop in performance.

### v2

@@ 5.5 Saliency Maps
To better understand the influence of the input features on our final predictions and also the difference our proposed method makes on how the networks determine what features to prioritize for prediction, we utilize a saliency map technique (Simonyan et al. (2013)) to visualize the importance of model weights on our input time axis and features. According to the predicted class $c$, for example, the decision-making process of the model can be represented as $S_{c}(I)=w_{c}^{T}I+b_{c}$, where $I,S_{c}(I),w,b$ are the model input, output, weights, and bias, respectively. The purpose of this method is to calculate the model weights on the input layer by gradients, for example, $w=\frac{\partial S_{c}}{\partial I}|_{I_{0}}$. The calculated $w$ represents the model saliency associated with the input layer. In this study, we fetch the saliency maps for all the test samples and calculate the average saliency map to develop the general intuition of important features and time steps.
We calculate the feature importance for our experiments according to the formulas in this section and plot them in the figures presented in Section 6.
@@ 6.1 ADULT Dataset
6.1.1 Fairness Analyses
The initial analysis before running a model shows that all 3 protected labels were biased by data. For example, for the age label, the DIR score is above 2 for both the training and testing data; for sex, it is approximately 0.36 for both, and for race, it is approximately 0.47 for both. After running the base model, this trend continues, where the model introduces even more bias by age, making the DIR above 3 and below 0.1 for sex, At the same time, for race, it stays similar to the initial analysis with a score of approximately 0.5.
6.1.2 Methods Implementation
Since our analyses on the baseline model show bias by all 3 chosen protected labels where their DIRs are beyond the acceptable range of 0.8-1.2, we experiment with our proposed method on all 3 protected labels. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figures 2. To give a better, more linear visual representation of our results, we calculate the absolute value of [$1-$DIR Score] for the models to plot against the accuracy. The lines on the plots represent the Pareto fronts which contain the models that perform best to choose from. For all of the experiments on this dataset, we opt for the models that provide us the best fairness each time due to the insignificant differences in performance (accuracy) between the models (they all fall between 0.77 and 0.8). Therefore, we choose the models enclosed in the green circles to represent our final models.
@@ 6.1.2 Methods Implementation
We compare our final performance results of the proposed method, reweighing, ARL, and FairRF in relation to the baseline performances as shown in Figure 3. In this figure, the x-axis represents the percentage improvement in the DIR score from the baseline in relation to the ideal score of 1, and the y-axis shows the percentage differences of the accuracy scores from the baseline accuracy. The reweighing technique improved the fairness scores, with the least impact on the accuracy of the models. For the protected age label, reweighing improved the DIR score by 84%, for sex by 15.9%, and for race by 00.67% which we deemed not significant. FairRF improved fairness for age and sex by approximately 45%, better than all other methods, but at a higher cost of performance. For the race label, FairRF diminished both fairness and accuracy of the model.
The chosen model from our proposed method performed better than reweighing in terms of fairness, with an improvement in the DIR score of the age label by 128%, sex by 32%, and 4.2% for the race label. However, this improvement in fairness came at a slight performance cost, where the proposed method caused an average decrease of 1.7% in accuracy for all 3 labels. The reweighing method had an even lower performance cost, improving the accuracy score for the sex label. Our implementation of ARL did not improve fairness of the sex and race labels, but significantly improved the DIR score for the age label by 114%, which is most comparable to the proposed method. This method also produced similar accuracy scores to the proposed method.
@@
6.1.3 Explainability
After calculating the feature importance for the baseline model and the models produced by our proposed method for age, race, and sex labels, the scores we obtained show us that our method redistributed weights to features less relevant to the protected label. This change is demonstrated in Figure 4. This figure visualizes the feature importance for the demographic labels as bar charts, which were chosen over saliency maps due to their clearer representation in scenarios without temporal data dependencies, making them easier to understand and interpret. The figure visualizes the feature importance for the demographic labels, and b represents the feature importance for the target label from the implementation of the base mode, and those of our proposed method for each of the protected labels. These figures show that the workclass feature holds the most weight across the board for all labels and remains equally important after the proposed method implementation. On the other hand, the relationship feature holds no weight in any of the prediction tasks across all experiments. The changes in importance are hence demonstrated more on the 3 remaining features: education, marital status, and occupation. Furthermore, the relative importance of these features changes across the experiments according to the demographic label we tested our proposed method on. For example, for the age label, the least important features are occupation, marital status, and education in that order, and the proposed method implemented with age shifted weight importance to these features. This trend is observed for the other labels of race and sex.
@@ 6.2 MIMIC-III Dataset
6.2.1 Fairness Analyses
The baseline model for IHM prediction using the MIMIC-III dataset results in Accuracy = 0.866, AUROC = 0.603, and AUPRC = 0.181. Our analysis of the MIMIC-III data shows that the baseline model is only biased by marital status out of the 4 protected labels we analyzed, with a DIR score of 1.3. Regarding age, gender, and insurance labels, the baseline model shows unsubstantial bias where the DIRs calculated for each fell within the acceptable range of 0.8-1.2, as shown in Table 3. The differences in FN scores are also significantly low across all the protected labels.
@@ 6.2.1 Fairness Analyses
6.2.2 Methods Implementation
Since our analyses show bias by marital status, we experiment with our proposed method using our target label of IHM and the protected label, marital status. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figure 5 for DIR plotted against the chosen metric AUROC. To provide a better, more linear visual representation of our results, we calculate the absolute value of [$1-$DIR Score] for the models to plot against the accuracy as we did in 6.1. The lines on the plots represent the Pareto fronts which contain the models that performed best to choose from for each of the performance scores. Our proposed method improves the 3 performance scores for IHM prediction across the 4 protected labels. The general accuracy for the models on the Pareto front shows an average improvement of 1.74% from the baseline model. The AUROC score improves by an average of 20%, and the AUPRC score by 19.7%. In this experiment, we focus on prioritizing the AUROC score, as mentioned in Section 5.2 because AUROC is prioritized for this dataset in previous studies (Harutyunyan et al. (2019)), and the accuracy variability for the models in the Pareto front is low (approximately 0.65%). We select our final model from the AUROC vs. DI score Pareto plot (Figure 5). From this plot, the 2 circled models represent the optimal solutions from our method. We choose model 40 as our final solution since it provided an almost equal improvement in DIR as model 23, with a more significant improvement in AUROC, AUPRC, and accuracy.
Figure 5: MIMIC-III: Pareto Front Plot for AUROC vs DI Ratio Score
Implementing the reweighing bias mitigation technique shows similar prediction performance when compared to our proposed method (as seen in Table 4), improving the AUROC, AUPRC, and accuracy slightly less than the proposed method when compared to the baseline. Reweighing improves the DIR score of the baseline model by 12% but does not perform as well as our proposed method. Regarding the difference in FN scores, the reweighing technique performs worse than the baseline. Implementation of ARL and FairRF performed comparably to the proposed method in terms of fairness considering DIR scores, but they diminished performance at significantly higher rates. There was an average drop of 7.9% in AUROC between the 2 methods.
@@ 6.2.2 Methods Implementation
Method Acc AUROC AUPRC DIR Diff in FN
Base Model 0.886 0.603 0.181 1.308 0.005
Reweighing 0.883 0.777 0.332 1.188 -0.05
ARL 0.719 0.528 0.136 0.954 0.064
FairRF 0.669 0.520 0.118 0.982 -0.035
Proposed Method 0.883 0.797 0.367 1.043 0.005
6.2.3 Explainability
The saliency maps from experiments on the MIMIC-III dataset can be found in Appendix C. The deeper red color in these figures represents features that hold more weight and, therefore, are more critical for target prediction. Evaluating the map for the MIMIC-III baseline model gives us an insight into some of the features that carried the most weight for IHM prediction. For example, the Glascow coma scale (GCS) motor response for Abnormal flexion and extension appears to be highly correlated with IHM prediction. GCS is a tool used to assess a patient’s level of consciousness after a brain injury (Jain and Iverson (2018)). The GCS motor response score evaluates the patient’s ability to obey commands or move in response to a stimulus. Numerous studies have examined the relationship between the GCS motor response score and in-hospital mortality (Yumoto et al. (2019); Sadaka et al. (2012); Settervall et al. (2011)). Generally, lower motor response scores are associated with a higher risk of mortality which explains the relevance of these features in our IHM prediction model.
Two other features that are shown to be highly correlated with the baseline IHM prediction are Fraction inspired oxygen (FiO2), which is a measure of the concentration of oxygen that a patient is receiving, usually through supplemental oxygen therapy (Fuentes and Chowdhury (2020)), and pH, which is a measure of the acidity or alkalinity of a patient’s blood (Arias-Oliveras (2016)). Evidence suggests that higher FiO2 levels may be associated with an increased risk of in-hospital mortality (de Jonge et al. (2008)), where higher FiO2 levels were associated with an increased risk of mortality in both adult and pediatric patients. There is also some evidence from prior studies that indicate an association between abnormal pH levels and an increased risk of in-hospital mortality (Park et al. (2020a)), where both low and high pH levels were associated with increased mortality in critically ill patients.
@@ 6.3 SNAPSHOT Dataset
6.3.1 Fairness Analyses
After performing the initial set of 32 fairness analyses on the SNAPSHOT data, our results show that the data itself does not contain substantial bias by any of the 8 protected labels we tested by. The scatter plot in Figure 6 shows that all scores fall within the acceptable range of 1.2 and 0.8 for the DIR score, except for the experiment with Race as a protected label and evening-sad-happy as the target label (specified by the red circle on the figure). For the 4 target labels, morning-sad-happy, morning-stressed-calm, evening-sad-happy, and evening-stressed calm, the F1 performance scores fell from approximately 0.7 to 0.8.
@@ 6.3.1 Fairness Analyses
6.3.2 Methods Implementation
We run our proposed method on Race as a protected label and evening-sad-happy as the target label and obtain accuracy scores in the range of 0.76 to 0.77 and DIR scores between 0.80 and 0.88 for the model ensemble produced. We choose the model with the best DIR score of 0.882 as shown in Figure 7. This is an improvement of approximately 0.1 in fairness from the baseline model with a DIR score of 0.78. In terms of performance, this method did not affect the F1 score when compared to the baseline model. Both have F1 scores of approximately 0.76. Using the reweighing technique, we obtain a perfect DIR score of 1, but that comes at a high cost of performance, with the F1 score dropping to 0.39, a loss of 37% in performance.
@@ 6.3.3 Data Manipulation on SNAPSHOT
After analyzing the SNAPSHOT baseline model for bias and our results showing that the data is only biased by the Race label for evening-sad-happy label prediction; we perform some data manipulation to create a more significant imbalance for the gender label in order to determine if this would decrease the fairness performance and to test our proposed method on it. Details on this process can be found in Appendix B.
After running the base model for all 4 target labels on this newly manipulated data and analyzing fairness for all 3 experiment settings, we obtain similar performance scores. The fairness scores for experiments S-1 and S-2 are similar, with only the DIR score of the target label of evening-sad-happy occurring outside of the acceptable bounds of 0.8-1.2. In contrast, the fairness scores for experiment S-3 did not change significantly from the original SNAPSHOT experiment with no data manipulation. The F1 and fairness metric scores for experiment S-2 are listed in Table 5. As seen in the table, the DIR score of the target label of evening-sad-happy (the gray highlighted cell) falls outside the acceptable bounds of 0.8 and 1.2. All other fairness scores fall within these acceptable bounds.
@@
After running our proposed bias mitigation algorithm on the label combination of evening-sad-happy and gender on the manipulated data, we obtain an ensemble of models whose performance and DIR scores are plotted in Figure 8. On average, the models have an F1 score of 0.74 which is a 5.8% drop in performance from the baseline evening-sad-happy model. The average DIR score of the 9 models on the Pareto front is 1.034, which is well within the acceptable bounds of 0.8-1.2 and an overall improvement of 0.257 from the base DI ratio score of 1.291. We chose models 89 and 43 (green and red circles in Figure 8) as potential final models. Model 89 has the best DI ratio score of 1.001, which is almost the ideal score of 1, but it comes with a diminished F1 score of 0.722, a 7.6% drop in performance. It differs in FP and FN rates scores of -0.082 and -0.084, respectively. Model 43 has a DI ratio score of 0.995, with slightly better performance (0.733) than model 89.
@@
6.3.4 Explainability
The saliency maps from experiments on the SNAPSHOT dataset can be found in Appendix C. The deeper red color in these figures represents features that hold more weight and are, therefore, more important for target prediction across the 7 timesteps. From the saliency maps for the baseline method on SNAPSHOT, a few features show the highest weights in the prediction task, at the top of this list of features being no_sleep_24, phys_3H-10H-medStepsWeightedPeaks, SMS_0H-3H_total_num. Comparing these maps to those from the model produced by our proposed method on SNAPSHOT, it can be noted that our proposed method distributed the weights more evenly among all the features in general, explicitly moving some of the importance away from the more highly-weighted features in the baseline model prediction.
@@ 7.1 Fairness and Performance
Facing the recurring challenges of bias and fairness in machine learning, our study introduces a bias mitigation technique that allows for tuning the trade-off between fairness and performance. Our method demonstrates enhanced fairness across three datasets—ADULT, MIMIC-III, and SNAPSHOT with minimal impact on performance and notable improvements for MIMIC-III. When compared to the standard reweighing bias mitigation technique, our approach shows significant improvements in fairness for the ADULT and MIMIC-III datasets and a slight improvement for SNAPSHOT, all without sacrificing performance. In contrast, reweighing reduces the F1 score by 37%. Overall, our method outperforms other state-of-the-art methods implemented on ADULT and MIMIC-III datasets in boosting fairness while maintaining performance.
@@
From our experiments in predicting income with the ADULT dataset, our analyses show us that the baseline model prediction of income was biased based on age, sex, and race, which corresponds with analyses from prior works (Le Quy et al. (2022); Lahoti et al. (2020); Friedler et al. (2019); Zafar et al. (2017)). Implementing our proposed method provides a distribution of models that improves the DIR scores to within the ideal threshold of 0.8 and 1.2. This method performs better than the reweighing technique in terms of fairness, leading to an average of 1.7% decrease in accuracy for all three protected labels.
The LSTM model predicting in-hospital mortality (IHM) using the MIMIC-III dataset exhibits bias by marital status but not by age, gender, or insurance type. This finding aligns partially with the fairness analysis by Röösli et al. (2022), which noted no gender bias. Unlike their study, which did not consider marital status, we include it based on evidence suggesting its significant health impact. Studies have shown that unmarried individuals face higher rates of adverse cardiovascular events and mortality compared to married ones (Dhindsa et al. (2020); Johnson et al. (2000)).
Upon identifying marital status bias, we apply our proposed bias mitigation technique, enhancing fairness to nearly optimal levels with a DIR of 1.04 and substantially equalizing FN rates across groups. Additionally, our method boosts model performance, increasing accuracy by 1.7%, AUROC by 19.4%, and AUPRC by 18.6%. This surpasses the performance improvements of the reweighing technique. Our findings support research suggesting that does not always lead to a drop in utility. Instead, it could lead to enhanced performance under specific circumstances (Deho et al. (2022)).
The ARL method shows promising results in improving fairness based on the age label for the ADULT dataset, but it falls short in addressing fairness concerns related to sex and race labels. This discrepancy could be linked to challenges similar to those observed with datasets like COMPAS, where ARL struggles to effectively address race-related fairness issues in the original publication. This indicates potential gaps in the ARL approach where race or other protected attributes are not computationally identifiable. Notably, the publication does not provide a detailed explanation for the method’s under-performance on certain datasets and attributes.
The FairRF method, developed by Zhao et al., introduces an approach that utilizes related features to develop fair classifiers without direct reliance on sensitive attributes. This method significantly enhances fairness for age and sex labels on the ADULT dataset by approximately 45%, surpassing other methods but at a considerable performance cost. However, for the ADULT race label, it reduces both fairness and accuracy, while it improves fairness for marital status on the MIMIC-III dataset, also at a significant performance cost. A critical limitation of the FairRF method is its heavy dependence on the selection and quality of related features. If the set of related features is incomplete or comprises noisy data, the effectiveness of the FairRFmethod may be significantly reduced. The success of this approach is highly dependent on how accurately the related features can represent the underlying sensitive attributes without introducing new biases.
@@
For SNAPSHOT, we analyze baseline models that predict 4 self-reported mood-related labels to determine if they are biased by 3 demographic labels, namely gender, race, and ethnicity, and 5 personality-based labels, namely openness, conscientiousness, extraversion, agreeableness, and neuroticism. These analyses determine that the DIR scores from all our experiments are within the chosen acceptable range of 0.8 to 1.2, except for the model predicting evening-happy-sad when analyzed by race. Implementing our proposed method improves the DIR score to 0.88, an improvement of 0.1 from that of the baseline method. In terms of performance, our proposed method has little effect in this case. Implementing the reweighing technique provides a perfect DIR score of 1 but a considerable performance cost of 37%.
@@
As our initial analyses on the SNAPSHOT dataset baseline models do not show substantial bias by the 8 protected labels, we manipulate the data by removing samples from 75 participants identified as female from the training data to determine the performance of our proposed method further. After running the baseline model on this altered data and analyzing, it shows bias by gender for the evening-sad-happy label with a DIR score of 1.29, slightly above our chosen acceptable range. Running our proposed technique results in an average, almost ideal DIR score of 1.034, an average 5.8% drop in performance.